How many molecules of hydrochloric acid will react with 5.2 grams of aluminum hydroxide?
1
Expert's answer
2013-02-28T11:22:32-0500
The chemical equation for thi reaction is
Al(OH)3 + 3HCl = AlCl3 + 3H2O
MW (HCl) = MW (H) + MW (Cl) = 1.0 + 35.5 = 36.5 g/mol
MW ( Al(OH)3 ) = MW (Al) + 3 x MW (O) + 3 x MW (H) = 27.0 + 3 x 16.0 + 3 x 1.0 = 78.0 g/mol

The amount of aluminium hydroxide

n ( Al(OH)3 ) = m ( Al(OH)3 ) / MW ( Al(OH)3 ) = 5.2 / 78.0 = 0.067 moles

The amount of hydrochloric acid

n (HCl) = n (Al(OH)3) x 3 = 0.067 x 3 = 0.201 moles

According to Avogadro law: & n(moles) = N (number of molecules) / Na

The number of molecules of HCl is

N = n x Na = 0.201 x 6.02 x 10^23 = 1.21 x 10^23 (molecules)

Answer:& N = 1.21 x 10^23 molecules of HCl